OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (cwrw238):

let the numbers be x x + 1 and x = 2 then x + x +1 + x + 2 = 267 now solve this for x x +2 is your largest number

OpenStudy (cwrw238):

can you do that?

OpenStudy (cwrw238):

3x + 3 = 267 whats x?

OpenStudy (anonymous):

i can do from her ty

OpenStudy (cwrw238):

yw